how to determine the slope and y-intercept of 2x-2y=6 and 3x+4y=12
Answer by stanbon(75887)

determine the slope and y-intercept of 2x-2y=6 and 3x+4y=12
------------
2x-2y=6
Solve for y to get:
x-y=3
y=x-3
slope = 1; y-int = -3
------------
3x+4y=12
y=(-3/4)x+3
slope = -3/4; y-int = 3
==================
Cheers,
Stan H.